OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form. –2x – y = –7 2x – y = –7 2x – y = 11 –2x + y = –11

OpenStudy (ivettef365):

you have y=mx+b 5 = (2)(-1) + b, so b = 7, so the equation would be y=2x+7

OpenStudy (anonymous):

okay and what's next

OpenStudy (ivettef365):

now in standard form, you can see which one will convert to y= 2x +7 by isolating the y, you know that the y-int =7 and not 11, so it must be between a or b

OpenStudy (anonymous):

is it A?

jimthompson5910 (jim_thompson5910):

if it was A, then plugging (-1, 5) into choice A should give a true statement –2x – y = –7 –2(-1) – (5) = –7 ... plug in x = -1 and y = 5 2 - 5 = -7 -3 = -7 but the last equation is false, so –2x – y = –7 is false when x = -1 and y = 5 So (-1, 5) is NOT a solution to –2x – y = –7

OpenStudy (anonymous):

so it is B

OpenStudy (ivettef365):

lets try a to see what we get -2x -y = -7 -y = 2x -7 --- -- -- -1 -1 -1 = y = -2x +7 what do you think

OpenStudy (anonymous):

the answer is B, right?

jimthompson5910 (jim_thompson5910):

let's try B 2x – y = –7 2(-1) – (5) = –7 -2 - 5 = -7 -7 = -7 that's true, so (-1,5) is indeed a solution to 2x - y = -7

jimthompson5910 (jim_thompson5910):

to confirm the slope for B is 2, simply solve for y and get it into y = mx+b form

OpenStudy (anonymous):

thanks guys so the answer is 2x-y= -7

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (ivettef365):

correct